OpenStudy (anonymous):

will medal & fan(: An equation is shown below: 5x + 3(x – 5) = 6x + 2x – 15 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation. (4 points)

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

@aloud

OpenStudy (anonymous):

I can solve the equation for you, as for the property I'm not sure 5x + 3(x - 5) = 6x + 2x - 15 5x + 3x - 15 = 8x - 15 8x - 15 = 8x - 15 8x = 8x x = x

OpenStudy (anonymous):

thank you so much i apreciate it! @Johnvdb

OpenStudy (anonymous):

No problem :)

OpenStudy (johnweldon1993):

Well @Johnvdb the first thing you did was "distribute" the 3 into the parenthesis.. so 1 property would be the distributive property!

OpenStudy (anonymous):

I woke up 5 minutes ago and logged on, so not all the information is going into my brain but thank you lol

OpenStudy (anonymous):

thanks(; @johnweldon1993

OpenStudy (johnweldon1993):

Lol no problem ;P